OpenStudy (anonymous):

A pair of parametric equations is given. x=sec(t), y=tan^2(t), 0 <= t <= pi/2 Find a rectangular-coordinate equation for the curve by eliminating the parameter.

OpenStudy (anonymous):

x2=1+y

OpenStudy (experimentx):

\[ y = \tan^2t = \sec^2t - 1 = x^2 - 1\] \[ y+1 = x^2 \]

OpenStudy (anonymous):

x belonging to [0,inf)